OpenStudy (anonymous):

what is the measure of angle T? angle s=122 degrees angle r=x degrees angle v= 90 degrees angle t = (x + 4) degrees. A. 68 degrees B.72 degrees C.76 degrees D.80 degrees

OpenStudy (anonymous):

R + S + T + V = 360 iff x + 122 + (x + 4) + 90 = 360 iff 2x + 216 = 360 iff 2x = 144 iff x = 72 T = x + 4 T = 72 + 4 = 76 degrees

OpenStudy (anonymous):

Ohhhh........ Now I get it thank u abayonmi12